With the following equation derive the various multiplier

Yd=Y-T+R

I=Io

G=Go

C=Co

Derive and compare the expenditure and transfer multipliers

2) supposed there is desrease of 300 million in government expenditures what effect will this have,no national income is give and the marginal propensity to consume is0.25
























1
Expert's answer
2021-03-09T15:45:12-0500

(1)

k=Y/(Y-C)

k=Y/(Y-I)

k=Y/(Y-G)

(2)

"300\/(1-0.25)=400"

Reduction of government expenditure reduces the aggregate demand.